OpenStudy (anonymous):

what is 8\45 / 1\63 compute in simplest form

OpenStudy (anonymous):

\[\frac{8}{45}\div \frac{1}{63}=\frac{8}{45}\times \frac{63}{1}\]

OpenStudy (anonymous):

cancel a 9 top and bottom get \[\frac{8\times 7}{5}\]